# Classification of locality preserving symmetries on spin chains

### v1

@@ Appendix F Proof of the main Theorem
Let $R$ be the range of $\alpha$ and take $L=16R$. By repeated use of Lemma E.6 we can stack with a decoupled symmetry $\alpha_{\operatorname{\mathrm{triv}}}=\bigotimes_{j\in L{\mathbb{Z}}}\,\rho_{j}$, where $\rho_{j}$ acts on site $j$ only, to obtain the symmetry $\alpha^{\prime}=\alpha\otimes\alpha_{\operatorname{\mathrm{triv}}}\sim\alpha$ of range $R$ on a stacked spin chain ${\mathcal{A}}^{\prime}$ such that $\alpha^{\prime}$ admits covariant right restrictions $\alpha_{\geq j}^{\prime}$ of defect size $7R$ at all sites $j\in L{\mathbb{Z}}$ and such that $g\mapsto\alpha_{\geq j}^{\prime(g)}$ are group homomorphisms.
@@
$\alpha_{\geq i}^{\prime(k)}\circ\alpha_{\geq j}^{\prime(\bar{k}gk)}\circ\alpha_{\geq i}^{\prime(\bar{k})}=\alpha^{\prime(k)}\circ\alpha_{\geq j}^{\prime(\bar{k}gk)}\circ\alpha^{\prime(\bar{k})}=\alpha_{\geq j}^{\prime(g)}.$ (F.1)
For each $j\in{\mathbb{Z}}$, define $\alpha_{j}^{\prime(g)}:=\alpha_{\geq jL}^{\prime(g)}\circ(\alpha_{\geq(j+1)L}^{\prime(g)})^{-1}$. Then it follows from Eq. (F.1) that
@@
    $[\alpha_{i}^{\prime(g)},\alpha_{j}^{\prime(h)}]=0$ whenever $i\neq j$.
@@
    $g\mapsto\alpha_{j}^{\prime(g)}$ are group homomorphisms.
@@
    $\alpha^{\prime}=\prod_{j\in{\mathbb{Z}}}\alpha^{\prime}_{j}$.
  4. 4.
    $\alpha^{\prime}_{j}$ is supported on $[jL-7R,(j+1)L+8R]$.
Pick unitaries $U_{j}(g)\in{\mathcal{A}}^{\prime}$ such that $\alpha_{j}^{\prime(g)}=\mathrm{Ad}[U_{j}(g)]$. Since $g\mapsto\alpha_{j}^{\prime(g)}$ is a group homomorphism, the unitaries $U_{j}(g)$ form a projective representation of $G$. The unitaries $U_{j}(g)$ commute with unitaries $U_{j^{\prime}}(h)$ whenever $\left|j^{\prime}-j\right|\geq 2$ because they have disjoint supports. Since $\alpha_{j-1}^{\prime(g)}$ commutes with $\alpha_{j}^{\prime(h)}$ we find that there are $\chi_{j}:G^{2}\rightarrow U(1)$ such that
@@
Our first goal is now to upgrade these unitaries $U_{j}$ to unitaries $\widetilde{U}_{j}$ with the same spatial support, and such that $g\mapsto\widetilde{U}_{j}(g)$ are still projective representations, but also such that
@@
In order to achieve this, we need to stack again. At each site $jL$ with $j\in{\mathbb{Z}}$, we add a degree of freedom with algebra ${\mathcal{C}}_{j}$ which is a copy of ${\mathcal{A}}_{[L(j-1),L(j+1)]}$, obtaining the enlarged spin chain ${\mathcal{A}}^{\prime\prime}$. We recall that the projective representations $U_{j-1}$ and $U_{j}$ map in ${\mathcal{A}}_{[L(j-1),L(j+1)]}$. Let us now consider the conjugate representations $\overline{U}_{j-1}$ and $\overline{U}_{j}$ mapping in the copy ${\mathcal{C}}_{j}$. The conjugate representations will always be considered inside ${\mathcal{C}}_{j}$ whereas the original representations act on the spin chain ${\mathcal{A}}^{\prime}$. We now define the symmetry $\rho_{Lj}^{\prime\prime}$ on the added degrees of freedom in ${\mathcal{C}}_{j}$ given by
@@
We note that $\widetilde{U}_{j}$ is a projective representation since it is a tensor product of projective representations. Writing $\alpha_{j}^{\prime\prime(g)}:=\mathrm{Ad}[\widetilde{U}_{j}(g)]$, we can verify that all properties of $\alpha_{j}^{\prime}$ listed below (F.1) hold for $\alpha_{j}^{\prime\prime}$ as well, in particular, $\alpha^{\prime\prime}=\bigotimes_{j\in{\mathbb{Z}}}\alpha_{j}^{\prime\prime}$. However, as announced, we now have the stronger commutation property (F.3). This is checked by an explicit computation using (F.2), (F.4) and $\chi_{j}(g,h)\bar{\chi}_{j}(g,h)=1$.
We now proceed to decouple $\alpha^{\prime\prime}$. This requires more stacking. For each $j\in{\mathbb{Z}}$ let $m(j)=jL+8R$ be the site sitting half way between $jL$ and $(j+1)L$. Stack with a spin chain ${\mathcal{B}}$ with on-site algebras ${\mathcal{B}}_{m(j)}=\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$ for all $j$ and ${\mathcal{B}}_{j}\simeq{\mathbb{C}}$ at all other sites to obtain a new spin chain ${\mathcal{A}}^{\prime\prime\prime}$. Let $\rho^{\prime\prime\prime}_{\operatorname{\mathrm{reg}}}$ be the decoupled symmetry on ${\mathcal{B}}$ which acts with the regular representation $\rho^{\prime\prime\prime}_{\operatorname{\mathrm{reg}}}$ on each ${\mathcal{B}}_{m(j)}$. Let $\alpha^{\prime\prime\prime}=\alpha^{\prime\prime}\otimes\rho^{\prime\prime\prime}_{\operatorname{\mathrm{reg}}}\sim\alpha^{\prime\prime}\sim\alpha$. We now explicitly construct a FDQC $\gamma$ such that $\gamma^{-1}\circ\alpha^{\prime\prime\prime}\circ\gamma$ is block partitioned.
Define unitaries
@@
$W_{j}(g):=V_{j}^{*}\,\big{(}\widetilde{U}_{j}(g)\otimes L_{m(j)}^{(g)}\big{)}\,V_{j}=\sum_{h\in G}\widetilde{U}_{j}(gh)^{*}\widetilde{U}_{j}(g)\widetilde{U}_{j}(h)\otimes|gh\rangle\langle g|=\sum_{h\in G}\,\operatorname{\mathbbm{1}}\,\otimes\,c_{j}(g,h)|gh\rangle\langle h|,$
for some 2-cocycle $c_{j}:G^{2}\rightarrow U(1)$. Here $L_{m(j)}^{(g)}\in{\mathcal{B}}_{m(j)}$ is the left regular action. The unitary $W_{j}(g)$ is seen to act non-trivially only on ${\mathcal{B}}_{m(j)}$ and so
@@
Let $\alpha$ and $\beta$ be symmetries on spin chains ${\mathcal{A}}$ and ${\mathcal{B}}$ respectively. If $\alpha$ and $\beta$ have the same anomaly then they are stably equivalent.
@@
We give the necessary definitions of (twisted) group cohomology and the slant product. For an in depth treatment, see for example the monograph [23].
Let $G$ be a group and $M$ a $G$-module. Consider the co-chain complex
@@
where $C^{n}$ is the collection of all functions $G^{n}\to M$ and $\delta$ is the differential map which sends a function $C^{n}\ni\theta:G^{n}\to M$ to $\delta\theta\in C^{n+1}$ defined by
$(\delta\theta)(g_{1},...,g_{n+1})=g_{1}\cdot\theta(g_{2},...,g_{n+1})\left(\prod\limits_{j=1}^{n}\theta(g_{1},...,g_{i}g_{i+1},...,g_{n+1})^{(-1)^{j}}\right)\theta(g_{1},...,g_{n})^{(-1)^{n+1}}.$
The map $\delta$ is called the co-boundary map, it satisfies $\delta^{2}=\mathrm{id}$.
The $n$-cocycles $Z^{n}$ are the functions in $C^{n}$ in the kernel of the co-boundary map $\delta$. The $n$-coboundaries $B^{n}$ are the functions of $C^{n}$ in the image of $\delta:C^{n-1}\to C^{n}$. The $n$-th cohomology group $H^{n}(G,M)$ is defined by $H^{n}(G,M)=Z^{n}/B^{n}$ together with the multiplication $[\theta]\cdot[\phi]=[\theta\phi]$ where $(\theta\phi)(g_{1},...,g_{n})=\theta(g_{1},...,g_{n})\cdot\phi(g_{1},...,g_{n})$ for two representatives $\theta,\phi\in Z^{n}$. The quotient is well defined because $\delta^{2}=1$.
In the following, and throughout the rest of this paper, the words cocycle, coboundary and cohomology will always refer to the case $M=U(1)$ with the trivial action of $G$ on $U(1)$. The words twisted cocycle, twisted coboundary and twisted cohomology will always refer to the case $M=U(1)[G]$, the collection of functions $G\to U(1)$ with the following action. For $C^{n}\ni\theta$ write $\theta_{h}$ for the assignment $\theta:G^{n}\to U(1)[G]:(g_{1},...,g_{n})\mapsto(h\mapsto\theta_{h}(g_{1},...,g_{n}))$. The action of $f\in G$ on $\theta$ is defined by
$(f\cdot\theta)(g_{1},...,g_{n})=(h\mapsto\theta_{\bar{f}hf}(g_{1},...,g_{n}))$
or pointwise by $f\cdot\theta_{h}=\theta_{\bar{f}hf}$.
A map $\omega:G^{3}\to U(1)$ is a (normal) 3-cocycle if
$\delta\omega(g_{1},g_{2},g_{3},g_{4})=\frac{\omega(g_{2},g_{3},g_{4})\omega(g_{1},g_{2}g_{3},g_{4})\omega(g_{1},g_{2},g_{3})}{\omega(g_{1}g_{2},g_{3},g_{4}),\omega(g_{1},g_{2},g_{3}g_{4})}=1.$ (G.1)

### v2

@@ F.1 Injectivity of $\Omega$
Let $R$ be the range of $\alpha$ and take $L=16R$. We stack ${\mathcal{A}}$ with local degrees of freedom $\operatorname{\mathrm{End}}(\mathbb{C}^{|G|})^{\otimes 3}$ at every site $jL$ for $j\in{\mathbb{Z}}$, obtaining a spin chain ${\mathcal{A}}^{\prime}$. Then, by Lemma E.9, there is a symmetry $\alpha^{\prime}\sim\alpha$ of range $R$ on the enlarged spin chain ${\mathcal{A}}^{\prime}$ that admits covariant right restrictions $\alpha_{\geq jL}^{\prime}$ of defect size $7R$ at sites $jL$ for all $j\in{\mathbb{Z}}$, and such that $g\mapsto\alpha_{\geq jL}^{\prime(g)}$ are group homomorphisms.
@@
$\alpha_{\geq iL}^{\prime(k)}\circ\alpha_{\geq jL}^{\prime(\bar{k}gk)}\circ\alpha_{\geq iL}^{\prime(\bar{k})}=\alpha^{\prime(k)}\circ\alpha_{\geq jL}^{\prime(\bar{k}gk)}\circ\alpha^{\prime(\bar{k})}=\alpha_{\geq jL}^{\prime(g)}.$ (F.2)
For each $j\in{\mathbb{Z}}$, define $\alpha_{j}^{\prime(g)}:=\alpha_{\geq jL}^{\prime(g)}\circ(\alpha_{\geq(j+1)L}^{\prime(g)})^{-1}$. Then
@@
    Eq. (F.2) implies that $g\mapsto\alpha^{\prime(g)}_{j}$ are group homomorphisms,
@@
    by definition, each $\alpha_{j}^{\prime}$ is supported on $[jL-7R,(j+1)L+8R]$,
@@
    covariance also implies that $[\alpha_{i}^{\prime(g)},\alpha_{j}^{\prime(h)}]=0$ whenever $i\neq j$.
Pick unitaries $U_{j}(g)\in{\mathcal{A}}^{\prime}$ such that $\alpha_{j}^{\prime(g)}=\mathrm{Ad}[U_{j}(g)]$. Since $g\mapsto\alpha_{j}^{\prime(g)}$ is a group homomorphism, the unitaries $U_{j}(g)$ form a projective representation of $G$. The unitaries $U_{j}(g)$ commute with unitaries $U_{j^{\prime}}(h)$ whenever $\left|j^{\prime}-j\right|\geq 2$ because they have disjoint supports (see figure 3).
[…]
Figure 3: The symmetry $\alpha^{\prime}$ can be seen as a conjugation by a FDQC $\prod\limits_{j\in\mathbb{Z}}U_{j}(g)$.
However, since $\alpha_{j-1}^{\prime(g)}$ commutes with $\alpha_{j}^{\prime(h)}$ we find that there are phases $\chi_{j}:G^{2}\rightarrow U(1)$ such that
@@
In order to obtain the statements of the lemma, our goal is now to upgrade these unitaries $U_{j}$ to unitaries $\widetilde{U}_{j}$ with the same spatial support, and such that $g\mapsto\widetilde{U}_{j}(g)$ are still projective representations, but also such that
@@
In order to achieve this, we need to stack again. At each site $jL$ with $j\in{\mathbb{Z}}$, we add a local degree of freedom ${\mathcal{C}}_{j}$ which is a copy of ${\mathcal{A}}^{\prime}_{[L(j-1),L(j+1)]}$, obtaining the enlarged spin chain ${\mathcal{A}}^{\prime\prime}$. We recall that the projective representations $U_{j-1}$ and $U_{j}$ map in ${\mathcal{A}}_{[L(j-1),L(j+1)]}$. Let us now consider the conjugate representations $\overline{U}_{j-1}$ and $\overline{U}_{j}$ mapping in the copy ${\mathcal{C}}_{j}$. The conjugate representations will always be considered inside ${\mathcal{C}}_{j}$ whereas the original representations act on the spin chain ${\mathcal{A}}^{\prime}$. We now define the symmetry $\rho_{Lj}^{\prime\prime}$ on the added algebras in ${\mathcal{C}}_{j}$ given by
@@
(see figure 4).
$-5L$$-4L$$-3L$$-2L$$-1L$$0L$$1L$$2L$$3L$$4L$$5L$$6L$$U_{-4}(g)$$U_{-2}(g)$$U_{0}(g)$$U_{2}(g)$$U_{4}(g)$$U_{-5}(g)$$U_{-3}(g)$$U_{-1}(g)$$U_{1}(g)$$U_{3}(g)$$U_{5}(g)$
$\overline{U}_{-5}(g)$
$\overline{U}_{-6}(g)$
$\overline{U}_{-4}(g)$
$\overline{U}_{-5}(g)$
$\overline{U}_{-3}(g)$
$\overline{U}_{-4}(g)$
$\overline{U}_{-2}(g)$
$\overline{U}_{-3}(g)$
$\overline{U}_{-1}(g)$
$\overline{U}_{-2}(g)$
$\overline{U}_{0}(g)$
$\overline{U}_{-1}(g)$
$\overline{U}_{1}(g)$
$\overline{U}_{0}(g)$
$\overline{U}_{2}(g)$
$\overline{U}_{1}(g)$
$\overline{U}_{3}(g)$
$\overline{U}_{2}(g)$
$\overline{U}_{4}(g)$
$\overline{U}_{3}(g)$
$\overline{U}_{5}(g)$
$\overline{U}_{4}(g)$
$\overline{U}_{6}(g)$
$\overline{U}_{5}(g)$
[…]
Figure 4: The unitaries $\tilde{U}_{j}(g)$ are defined as tensor products of $U_{j}(g)$ with projective representations $\overline{U}_{j}(g)$ defined on appropriate degrees of freedom $\mathcal{C}_{j},\mathcal{C}_{j+1}$.
We note that $\widetilde{U}_{j}$ is a projective representation since it is a tensor product of projective representations. Writing $\alpha_{j}^{\prime\prime(g)}:=\mathrm{Ad}[\widetilde{U}_{j}(g)]$, we can verify that all properties of $\alpha_{j}^{\prime}$ listed below (F.2) hold for $\alpha_{j}^{\prime\prime}$ as well, in particular, $\alpha^{\prime\prime}=\prod_{j\in{\mathbb{Z}}}\alpha_{j}^{\prime\prime}$. However, as announced, we now have the stronger commutation property (F.4). This is checked by an explicit computation using (F.3), (F.5) and $\chi_{j}(g,h)\bar{\chi}_{j}(g,h)=1$.
Proposition F.2.
Proof.
By Lemma F.1, the symmetry $\alpha$ is stably equivalent to a symmetry $\alpha^{\prime\prime}$ that splits into local components which commute with each other. We proceed in constructing an explicit equivalence between $\alpha^{\prime\prime}$ and a block partitioned QCA. Let $R$ denote the range of $\alpha$, and let $L=16R$, exactly as in the proof of the previous Lemma. We use freely the notation from Lemma F.1.
The decoupling of $\alpha^{\prime\prime}$ requires more stacking. For each $j\in{\mathbb{Z}}$ let $m(j)=jL+8R$ be the site sitting half way between $jL$ and $(j+1)L$. Stack with a spin chain ${\mathcal{B}}$ with on-site algebras ${\mathcal{B}}_{m(j)}=\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$ for all $j$ and ${\mathcal{B}}_{j^{\prime}}\simeq{\mathbb{C}}$ at all other sites $j^{\prime}$ to obtain a new spin chain ${\mathcal{A}}^{\prime\prime\prime}$. Let $\rho^{\prime\prime\prime}_{\operatorname{\mathrm{reg}}}$ be the decoupled symmetry on ${\mathcal{B}}$ which acts with the left regular representation $\rho_{\operatorname{\mathrm{reg}}}$ on each ${\mathcal{B}}_{m(j)}$. Let $\alpha^{\prime\prime\prime}=\alpha^{\prime\prime}\otimes\rho^{\prime\prime\prime}_{\operatorname{\mathrm{reg}}}\sim\alpha^{\prime\prime}\sim\alpha$. We now explicitly construct a FDQC $\gamma$ such that $\gamma^{-1}\circ\alpha^{\prime\prime\prime}\circ\gamma$ is block partitioned. Define unitaries
@@
$W_{j}(g):=V_{j}^{*}\,\big{(}\widetilde{U}_{j}(g)\otimes(\rho_{\operatorname{\mathrm{reg}}})_{m(j)}^{(g)}\big{)}\,V_{j}=\sum_{h\in G}\widetilde{U}_{j}(gh)^{*}\widetilde{U}_{j}(g)\widetilde{U}_{j}(h)\otimes|gh\rangle\langle g|=\sum_{h\in G}\,\operatorname{\mathbbm{1}}\,\otimes\,c_{j}(g,h)|gh\rangle\langle h|,$
for some 2-cocycle $c_{j}:G^{2}\rightarrow U(1)$. Here $(\rho_{\operatorname{\mathrm{reg}}})_{m(j)}^{(g)}\in{\mathcal{B}}_{m(j)}$ is the left regular action on site $m(j)$. The unitary $W_{j}(g)$ is seen to act non-trivially only on ${\mathcal{B}}_{m(j)}$ and so
@@
The map $\Omega:\operatorname{\mathsf{Sym}}_{G}\rightarrow H^{3}(G,U(1))$ is injective. In other words, let $\alpha$ and $\beta$ be symmetries on spin chains ${\mathcal{A}}$ and ${\mathcal{B}}$, respectively. If $\alpha$ and $\beta$ have the same anomaly, then they are stably equivalent.
@@ F.2 Classification of locality preserving symmetries
G.1 Group cohomology
We give the necessary definitions of (twisted) group cohomology and the slant product. For an in depth treatment, see for example the monograph [34]. Recall that for a group $G$, a $G$-module is an abelian group $M$ equipped with a left $G$-action __$\_\cdot\_\ :G\times M\longrightarrow M$ satisfying
  1. 1.
    $g\cdot(xy)=(g\cdot x)(g\cdot y)$ for all $x,y\in M$, $g\in G$;
  2. 2.
    $g\cdot 0=0$,
where we use multiplicative notation for the module operation. Let $G$ be a group and $M$ a $G$-module. Consider the cochain complex
@@ G.1 Group cohomology
where $C^{n}:=C^{n}(G,M)$ is the collection of all functions $G^{n}\to M$ and $\delta$ is the differential map which sends a function $C^{n}\ni\theta:G^{n}\to M$ to $\delta\theta\in C^{n+1}$, defined by
$(\delta\theta)(g_{1},...,g_{n+1})=g_{1}\cdot\theta(g_{2},...,g_{n+1})\left(\prod\limits_{j=1}^{n}\theta(g_{1},...,g_{j}g_{j+1},...,g_{n+1})^{(-1)^{j}}\right)\theta(g_{1},...,g_{n})^{(-1)^{n+1}}.$ (G.1)
The differential map $\delta$ is also called the coboundary map, and it satisfies $(\delta\circ\delta)(\theta)=1$. The $n$-cocycles $Z^{n}$ are the functions in $C^{n}$ in the kernel of the coboundary map $\delta$. The $n$-coboundaries $B^{n}:=B^{n}(G,M)$ are the functions of $C^{n}$ in the image of $\delta:C^{n-1}\to C^{n}$. The $n$-th cohomology group $H^{n}(G,M)$ is defined by $H^{n}(G,M)=Z^{n}/B^{n}$ together with the multiplication $[\theta][\phi]=[\theta\phi]$ where $(\theta\phi)(g_{1},...,g_{n})=\theta(g_{1},...,g_{n})\phi(g_{1},...,g_{n})$ for two representatives $\theta,\phi\in Z^{n}$. The quotient is well defined because of the property $(\delta\circ\delta)(\theta)=1$.
In the following, and throughout the rest of this paper, the words cocycle, coboundary and cohomology will always refer to the case $M=U(1)$ with the trivial action of $G$ on $U(1)$, namely, the action $g\cdot\theta=\theta$ for all $g\in G$.
G.1.1 Third group cohomology
Since the anomaly of a locality preserving symmetry takes values in degree three group cohomology, we explicitly state both the cocycle and coboundary conditions: A map $\omega:G^{3}\to U(1)$ is a (normal) 3-cocycle if
$(\delta\omega)(g_{1},g_{2},g_{3},g_{4})=\frac{\omega(g_{2},g_{3},g_{4})\omega(g_{1},g_{2}g_{3},g_{4})\omega(g_{1},g_{2},g_{3})}{\omega(g_{1}g_{2},g_{3},g_{4}),\omega(g_{1},g_{2},g_{3}g_{4})}=1.$ (G.2)
@@ G.1.1 Third group cohomology
G.2 Twisted group cohomology
The $G$-module $U(1)[G]$ consists of maps $\lambda:G\to U(1)$, assigning $g\mapsto\lambda_{g}$, with a left $G$-action
$(k\cdot\lambda)_{g}=\lambda_{\bar{k}gk},\qquad k,g\in G.$
This is a left action as $(k\cdot l\cdot\lambda)_{g}=(l\cdot\lambda)_{\bar{k}gk}=\lambda_{\overline{kl}gkl}=(kl\cdot\lambda)_{g}$. Abelian multiplication is given pointwise by $(\lambda\lambda^{\prime})_{g}:=\lambda_{g}\lambda^{\prime}_{g}$, where $\lambda_{g}\lambda^{\prime}_{g}$ is $U(1)$ multiplication. The module $U(1)[G]$ is a $G$-graded crossed module $\bigoplus_{g\in G}U(1)_{g},$ where each $U(1)_{g}$ is a copy of the unitary group $U(1)$ and $G$ acts as conjugation on the label, i.e. $k\cdot U(1)_{g}\subseteq U(1)_{kg\bar{k}}$ [35, 36].
Example G.1 (First degree twisted cohomology).
By applying the definition of the differential (G.1), we check that a map $c:G\to U(1)[G]$ is a twisted 1-cocycle if
$(\delta c)_{h}(g_{1},g_{2})=\frac{c_{\bar{g}_{1}hg_{1}}(g_{2})c_{h}(g_{1})}{c_{h}(g_{1}g_{2})}=1,\qquad h,g_{1},g_{2}\in G,$ (G.4)
and a twisted 1-coboundary is a map
$(\delta\eta)_{h}(g)=\frac{\eta_{\bar{g}hg}}{\eta_{h}},\qquad g,h\in G,$ (G.5)
for some $\eta\in U(1)[G]$.
Example G.2 (Second degree twisted cohomology).